The sum of two decimal numbers is 5.3. Their difference is 1.7, and their product is 6.3. what are the two numbers
Answer by ankor@dixie-net.com(22740)

The sum of two decimal numbers is 5.3. Their difference is 1.7, and their product is 6.3.
what are the two numbers
:
Write an eq for the sum and an eq for diff, add the two equations
x + y = 5.3
x - y = 1.7
------------adding eliminates y, find x
2x = 7.0
x = 7/2
x = 3.5
:
Find y
3.5 + y = 5.3
y = 5.3 - 3.5
y = 1.8
:
Confirm this by finding the product: 3.5*1.8 = 6.3